OpenStudy (anonymous):

Which is 63 × 64 in exponential form? A. 36^12 B. 7^6 C. 6^12 D. 6^7

OpenStudy (anonymous):

would you multiply 63 and 64 then try each choice?

OpenStudy (anonymous):

If you do that...well, that seemed the corect approach to me, I would multiply and group factor them to see how can I put it in a form of x^y...but none of the values that the choices give seems to be equal to 63 × 64 (which equals 4032).

OpenStudy (anonymous):

Perhaps you wrote the options or the question incorrectly?

OpenStudy (anonymous):

No, I copied and pasted it. that is why I put it on here. It confused me.

OpenStudy (anonymous):

Is this from an online assignment?

OpenStudy (anonymous):

ok then try a factor tree

OpenStudy (anonymous):

it is not really an assignment. it is optional for us to do. but with crct approaching I need to try to study study study

OpenStudy (anonymous):

jasime, the values are not equal, so factoring wouldn't help here.

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

64=2^6 63=3^2 ×7 but that doesn't help, I am just simplifying them, just like I would to simplify a square root of these numbers... @noneyabusiness , ask your teacher about this problem.

OpenStudy (anonymous):

Ok, I will. Thanks anyways!

OpenStudy (anonymous):

Question him/her saying something along the following lines... the values that the options multiply to are not equal to 63 × 64... Good luck with your assignment! You welcome!